Suppose the perfectly competitive firm’s cost function is TC=Q3-21Q2+600Q+1820 and its AR=600. (2Pts)                         

    a. What is the profit maximization level of output?

b.     What is the maximum of profit or minimum of loss?


1
Expert's answer
2022-01-02T18:16:01-0500

a. P = MR = MC

"(TC=Q3-21Q2+600Q+1820)'=3Q^2-42Q+600"

With perfect competition , AR = MR

"3Q^2-42Q+600=600"

"3Q^2-42Q=0"

"Q(3Q-42)=0"

Q=0

3Q-42=0

Q=14

b. In conditions of perfect competition, the marginal income is equal to the price of the goods. Therefore, the rule MR = MC can be presented in another form: P = MR = MC, or P = MS. That is, in conditions of perfect competition, profit maximization (loss minimization) is achieved when the volume of production corresponds to the point of equality of marginal costs and price.